OpenStudy (anonymous):

which number is the least common multible of 6 and 15? 60 30 15 or 3

OpenStudy (cwrw238):

30

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

general way to do this is to find prime factors of the numbers: 6 = 2*3 15 = 3*5 3 is common to both so only use it once LCM = 2*3*5 = 30

OpenStudy (anonymous):

Which fraction is equivalent to 3 over 7?